OpenStudy (anonymous):

sinx/1-cosx +sinx/1+cosx = 2cscx

OpenStudy (anonymous):

Combine the fractions by finding a common denominator.

OpenStudy (anonymous):

\[\frac{\sin x}{1-\cos x}+\frac{\sin x}{1+\cos x}=2\csc x,~\text{right?}\]

OpenStudy (anonymous):

$$\frac{\sin x}{1-\cos x}+\frac{\sin x}{1+\cos x}=\frac{(1+\cos x)\sin x+(1-\cos x)\sin x}{(1-\cos x)(1+\cos x)}$$

OpenStudy (anonymous):

It's just a matter of simplifying now...

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

i got 2sinx

OpenStudy (anonymous):

Nope, try again

OpenStudy (anonymous):

$$\frac2{\sin x}=2\frac1{\sin x}=2\csc x$$

OpenStudy (anonymous):

how do you get sinx as denomenator

mathslover (mathslover):

\(\sin^2 x + \cos^2 x = 1\) \(\sin^2 x = 1 - \cos^2 x\)

mathslover (mathslover):

You have : (1-cos x)(1+cos x) as denominator. Use : (a+b)(a-b) = a^2 - b^2 identity. You get : \((1-\cos x)(1 + \cos x) = 1- \cos^2 x\) As : \(1-\cos^2 x = \sin^2 x \) Therefore, you get denominator as : \(\sin^2 x\)

mathslover (mathslover):

In the numerator you had : \(2\sin x\) while in the denominator you have : \(\sin^2 x\) \(\cfrac{2\sin x}{\sin^2 x} \)

mathslover (mathslover):

Cancel the terms and simplify.

OpenStudy (anonymous):

$$(1-\cos x)(1+\cos x)=1-\cos^2x=\sin^2x$$ is what our denominator simplifies to

OpenStudy (anonymous):

$$(1+\cos x)\sin x+(1-\cos x)\sin x=(1+\cos x+1-\cos x)\sin x=2\sin x$$is our numerator

mathslover (mathslover):

^^

OpenStudy (anonymous):

$$\frac{2\sin x}{\sin^2 x}=\frac2{\sin x}=2\frac1{\sin x}=2\csc x$$